OpenStudy (anonymous):

Mia Kaminsky wants to attend Riverside Community College. She will need to have $25,000 six years from today. Mia is wondering what she will have to put back in the bank today so she will have $25,000 six years hence. Her bank pays 5 percent compounded semiannually. By using tables in the handbook, the amount Mia will have to deposit is: $18,950 $33,226 $18,590 $33,622 None of these

OpenStudy (anonymous):

The Compound Interest Equation P = C (1 + r/n)^nt where P = future value C = initial deposit r = interest rate (expressed as a fraction: eg. 0.06) n = # of times per year interest in compounded t = number of years invested 25000 = C (1 + 0.05/2)^(2x6) I assume you can solve for C?? Works out to somewhere between 18000 and 19000. If you need help with working out C, let me know.